# Berkeley, CA 94720, USA

### v2

@@ Abstract
The Witten index of space of the $(2,0)$-theory compactified on spaces of the form $S^{3}/\Gamma\times S^{2}$, with a freely acting group $\Gamma$, and with external string sources implemented via timelike surface operator insertions, is expressed in terms of Ray-Singer torsion of $S^{3}/\Gamma$ and characters of irreducible representations of $\Gamma$. We compute it explicitly for the Dicyclic groups $\Gamma=\text{Dic}_{k}$. The torsion and characters are generally irrational numbers, but they nicely combine to an integer index. The results suggest fusion relations among the string sources with integer coefficients expressed in terms of Chebyshev polynomials.
@@ 1 Introduction
We consider the $(2,0)$-theory Witten:1995zh associated with the Lie group SU$(2)$ ($A_{1}$) (i.e., the low energy limit of the worldvolume theory of two coincident M5 branes after the center of mass is decoupled Strominger:1995ac ) along a Lorentzian manifold ×$\mathbb{R}\times X_{5}$, with a topological twist to preserve supersymmetry, where $X_{5}$ is a generic 5-dimensional smooth Riemannian manifold with a positive scalar curvature. We will later set $X_{5}$ to a quotient of either $S^{5}$ or $S^{3}\times S^{2}$ by a freely acting finite group $\Gamma$. Compactifying on $S^{3}/\Gamma\times S^{2}$ leads to $\mathcal{N}=2$ supersymmetry in 0+1d, while compactifying on $S^{5}/\Gamma$ leads to $\mathcal{N}=1$.
There is an ongoing effort to propose a fundamental formulation of the $(2,0)$-theory, and we list several approaches that have been suggested. (This list is by no means comprehensive and extensive reviews of the $(2,0)$-theory can be found in Moore2012 and Heckman:2018jxk .) The matrix model formulation of the $(2,0)$-theory on spaces with an $\mathbb{R}^{1,1}$ factor, based on instanton quantum mechanics, was proposed in Aharony:1997th , following an approach closely related to the BFSS matrix model Banks:1996vh . Another approach, known as deconstruction, was introduced in Arkani-Hamed:2001wsh . The conjecture is that in the large $N$ limit, 4d $\mathcal{N}=2$ quiver gauge theories on the Higgs branch reproduce the $(2,0)$-theory.
A third proposal involves studying the $(2,0)$-theory compactified on an $S^{1}$ circle as a 5d SYM theory Douglas:2010iu ; Lambert:2010iw ; Lambert:2012qy . This idea has its origins in observations made in Seiberg:1997ax . Another approach involves the use of 3-algebras to obtain an action for multiple M5 branes, as discussed in Lambert:2010wm . This approach was motivated by the success of BLG models Bagger:2006sk ; Gustavsson:2007vu for multiple M2 branes. Finally, we mention a twistor approach to non-abelian higher gauge fields introduced in Saemann:2011nb .
Independently of the full UV description of the $(2,0)$-theory, the fundamental observation Seiberg:1997ax that 5d SYM describes the low-energy limit of the 6d theory compactified on $S^{1}$ was employed to extract remarkable exact results, using supersymmetric localization Kallen:2012cs ; Kim:2012ava ; Kim:2013nva , and it will be our main tool as well.
@@
Ccompactification on three manifolds Dimofte:2011ju results in 3d $\mathcal{N}=2$ theories referred to as Class $\mathcal{R}$. These theories exhibit a rich duality structure including mirror symmetry. The last compactification we would like to mention, in the increasing dimension of compactified space, is the compactification on four manifolds $S^{3}\times S^{1}$. This results in a zero area limit of 2d q-deformed SYM theory Gadde:2009kb , where the q-deformation introduces a quantum group structure for the gauge group. This theory can actually be used to compute superconformal indices of $\text{Class}\;\mathcal{S}$ theories Gadde:2010te ; Gadde:2011ik .
In this paper, we compute the Witten index Witten:1982df of space of ground states of the 6d $\mathcal{N}=(2,0)$ theory with SU$(2)$ gauge group, compactified on a few particular examples of $X_{5}$, with a supersymmetry-preserving R-symmetry twist. We carry out the calculations by compactifying the time direction into a Euclidean circle $S^{1}$ and shrinking the circle to small radius. In this limit Seiberg:1997ax , the 6d theory is described by an effective 5d SYM with gauge group SU$(2)$, and one finds that the circle’s radius is proportional to a square of the gauge coupling constant of the resulting 5d SYM theory. The R-symmetry twist becomes a topological twist. Hence, we can compute the Witten index by calculating the partition function of the 5d SYM theory on $X_{5}$. As explained in Aharony:1998qu ; Henningson:2010rc ; Tachikawa:2013hya , additional discrete data in the form of an element ${\mathfrak{v}}\in H^{3}(X_{5},\mathbb{Z}_{2})$ needs to be specified in order to uniquely specify what is meant by space of the $(2,0)$-theory on $X_{5}$, and in 5d SYM ${\mathfrak{v}}$ specifies which gauge $\text{SU}(2)/\mathbb{Z}_{2}$ gauge bundles are to be included in the path integral.
@@
Similarly to what was originally found for 4d SYM Witten:1994cg ; Vafa:1994tf , the partition function of the 5d SYM theory and the correlation functions of SUSY Wilson loops are independent of the coupling constant and the metric because of the topological twist, which in 5d inserts a background gauge field for the $\mathfrak{so}(5)_{R}$ R-symmetry algebra that matches the geometrical spin connection of $X_{5}$. It was shown in Bershadsky:1995qy that the worldvolume theories that describe the dynamics of branes wrapping analytic submanifolds of a Calabi-Yau space naturally become topologically twisted. In our case, we can view the topological twist as the result of wrapping M5-branes on the zero section of the 10d space $T^{\star}X_{5}$ (the cotangent space of $X_{5}$). The spatial slices of the brane will reside along $X_{5}\subset T^{\star}X_{5}$, and the respective topological twist identifies the $\mathfrak{so}(5)$ holonomy algebra with the $\mathfrak{so}(5)_{R}$ R-symmetry algebra, both are subalgebras of the original real superalgebra of the 6d $\mathcal{N}=(2,0)$ theory: $\mathfrak{osp}(8^{*}|2)$. This particular topological twist of the 5d SYM theory complexifies the gauge field and also results in an additional 5d Chern Simons theory with Grassmann odd sections. This twist was discussed first in Geyer:2002gy and then in Bak:2015hba ; Bak:2017jwd .
@@
In what follows, we choose $X_{5}$ to be $S^{3}/\Gamma\times S^{2}$. Here, $\Gamma$ is one of the finite subgroups of SU$(2)$ of A or D type, which are constructed by the ADE McKay correspondence mckay1981graphs . For A-type, $\Gamma$ is a cyclic group $\mathbb{Z}_{k}$, and $S^{3}/\Gamma$ is a Lens space denoted by $L(k,1)$. [^1] In general one can add integers $l_{1},\dots,l_{n}$ coprime with $k$, where $n$ is the complex dimension of orbifold but in this paper we set them all to 1. For D-type, $\Gamma$ is a dicyclic group $\text{Dic}_{k}$ of order $4k$, and the resulting orbifold is referred to as a prism manifold. We will also briefly discuss $S^{5}/\Gamma$ with $\Gamma$ a cyclic group $\mathbb{Z}_{k}$ of order $k$. We review these geometries and possible extensions in detail in the first part of this paper.
We are able to compute the partition function on these geometries by applying supersymmetric localization methods. The localization equations result in BPS solutions that are given by flat connections. The $X_{5}$ spaces that we consider will have non-trivial fundamental groups and consequently, there will be non-trivial flat connections. After the path integral localization, we will need to compute functional determinants of twisted Hodge Laplacians on orbifolds. We will find that we can take a computationally simpler approach by using the Ray-Singer torsion which allows us to extract the answer by simply studying the group theoretic properties of representations of $\Gamma$.
@@
The outline of this paper is as follows. In §2, we provide and review basic facts about the geometrical and topological properties of the manifolds we consider in this paper (for the role of $X_{5})$. In §3, we describe the topological twist and the emergence of fermionic Chern-Simons theory (FCS). We also review how the partition function of FCS can be computed using Ray-Singer torsion. In §4, we provide an explicit calculation of torsion for $\Gamma=\mathbb{Z}_{k}$ and $\Gamma=\text{Dic}_{k}$: in the latter case we derive a novel closed form expression. We use this calculation to compute and list the Witten index in §5 for a few values of $k$. We then introduce timelike surface operators into the Witten index and prove that the resulting correlation function is always an integer. We also construct recurrence relations for these indices that could be interpreted as fusion relations for external string sources. In section §6, we conclude our results with a discussion of possible extensions.
@@ 2 Geometry of Spherical Orbifolds
We begin with a discussion of the various manifolds $X_{5}$ on which we will study the (twisted) $A_{1}$ 6d $\mathcal{N}=(2,0)$ theory, and we review some important facts about their homology groups. Recall that after compactification on a small $S^{1}$, we get a twisted version of 5d Super Yang-Mills theory with gauge group $G=\text{SU}(2)$ on the manifold $X_{5}$, which is a topological field theory that we will review below. To compute the partition function for a topological field theory that arises in this compactification, we will need to add the contributions coming from non-trivial flat connections of the SU$(2)$ gauge field. To get a non-trivial result, we will need non-trivial flat connections, which require non-trivial elements of the fundamental group $\pi_{1}(X_{5})$. Therefore, we want to pick manifolds $X_{5}$ that have a non-trivial fundamental group $\pi_{1}(X_{5})$. We will need to calculate the Ray-Singer torsion of the manifolds and to obtain non-vanishing results we will need to address the issue of fermionic zero modes. Ideal candidates for $X_{5}$ are quotients of spheres or products of spheres by a finite subgroup $\Gamma$ of the spin group.
@@
For $S^{3}/\Gamma\times S^{2}$, we take $\Gamma$ to be a subgroup of the spin group $\text{Spin}(3)\times\text{Spin}(4)\simeq\text{SU}(2)_{1}\times\text{SU}(2)_{2}\times\text{SU}(2)_{3}$, where we added subscripts $1,2,3$ to distinguish the three SU$(2)$ factors, and where SU$(2)_{1}$ acts on $S^{2}$, and $\text{SU}(2)_{2}\times\text{SU}(2)_{3}$ acts on $S^{3}$. In general, one can consider two choices that act freely: either to embed $\Gamma$ as a (finite) subgroup in SU$(2)_{2}$ [or, equivalently, in SU$(2)_{3}$], or to embed as a diagonal subgroup in $\text{SU}(2)_{1}\times\text{SU}(2)_{2}$, thus acting non-trivially both on $S^{2}$ and $S^{3}$. Collectively we can denote both choices as $(S^{3}\times S^{2})/\Gamma$. In either case, $\Gamma$ will be associated with an A, D, or E type Dynkin diagram, by the McKay correspondence. In the present paper, we focus on the first case where $\Gamma$ acts non-trivially on $S^{3}$. We perform the calculations for two choices of $\Gamma$: $\mathbb{Z}_{k}$ and Dick, which by the McKay correspondence have Dynkin diagrams of $A_{k-1}$ and $D_{k+2}$ type. Both groups act freely, leading to a smooth $X_{5}$.
@@
$r^{2k}=1,\quad s^{2}=r^{k},\quad s^{-1}rs=r^{-1}.$ (4)
We construct irreducible representations $\text{Dic}_{k}\rightarrow\text{SU}(2)$ (see also previous work Ju:2023umb ; Albrychiewicz:2024fkr that used a similar construction), which satisfy these relations for each holonomy and are given by two-dimensional matrices with unit determinant. From (4) one observes that there are abelian and non-abelian representations, which makes the manifold quotient by the dicyclic group more interesting than a quotient by a cyclic group only. We analyze these representations separately for even and odd $k$.
$\begin{array}[]{|c|c|c|}\hline\cr\text{Representation reference}&g(r)&g(s)\\ \hline\cr{\mathbf{1}_{{{1}}}}&1&1\\ \hline\cr{\mathbf{1}_{{{2}}}}&1&-1\\ \hline\cr{\mathbf{1}_{{{3}}}}&-1&1\\ \hline\cr{\mathbf{1}_{{{4}}}}&-1&-1\\ \hline\cr\end{array}$
@@
For any even $k$, there are four one-dimensional representations, listed in Table 1 together with the notation we use to refer to them (${\mathbf{1}_{{{1}}}},\dots,{\mathbf{1}_{{{4}}}}$). Associated to them are four inequivalent embeddings of $\text{Dic}_{k}$ in SU$(2)$:
@@
where $m=1,2,\dots,k-1$, and we refer to the representation associated with a particular $m$ as ${\mathbf{2}_{{m}}}$. For odd $m$, both matrices have determinant $1$, and the representation forms an embedding of Dick in SU$(2)$.
For odd $k$, the abelian irreducible representations are listed in Table 2, and the inequivalent embeddings $\text{Dic}_{k}\rightarrow\text{SU}(2)$ are listed in Table 3; there are $3$ of them. As for the non-abelian representations, the only difference from (6) is, that for odd $k$, $m$ now runs up to $k-2$.
@@
$\left({\mathbf{2}_{{m}}}\otimes{\mathbf{2}_{{m}}}\right)_{S}=\left\{\begin{array}[]{ll}{\mathbf{2}_{{2m}}}\oplus{\mathbf{1}_{{{1}}}},&\;\text{for $m\in 2\mathbb{Z}$}\\ {\mathbf{2}_{{2m}}}\oplus{\mathbf{1}_{{{2}}}},&\;\text{for $m\in 2\mathbb{Z}+1$}\end{array}\right.$ (7)
$\begin{array}[]{|c|c|c|}\hline\cr\text{Representation name}&g(r)&g(s)\\ \hline\cr{\mathbf{1}_{{{1}}}}&1&1\\ \hline\cr{\mathbf{1}_{{{2}}}}&1&-1\\ \hline\cr{\mathbf{1}_{{{3}}}}&-1&i\\ \hline\cr{\overline{\mathbf{1}}_{{{3}}}}&-1&-i\\ \hline\cr\end{array}$
@@
$\begin{array}[]{|c|c|}\hline\cr\text{Representation name}&\text{Decomposition into irreducible representations}\\ \hline\cr{\widetilde{\mathbf{2}}_{{{1}}}}&{\mathbf{1}_{{{1}}}}\oplus{\mathbf{1}_{{{1}}}}\\ \hline\cr{\widetilde{\mathbf{2}}_{{{2}}}}&{\mathbf{1}_{{{2}}}}\oplus{\mathbf{1}_{{{2}}}}\\ \hline\cr{\widetilde{\mathbf{2}}_{{{3}}}}&{\mathbf{1}_{{{3}}}}\oplus{\overline{\mathbf{1}}_{{{3}}}}\\ \hline\cr\end{array}$
Table 3: For odd $k$, embeddings of Dick in SU$(2)$ are constructed as direct sums of pairs of abelian representations from Table 2.
@@
We begin with the case of $S^{5}$ which has a spin isometry group Spin$(6)$, which is isomorphic to SU$(4)$. The various possible choices of $\Gamma$ that act freely on $S^{5}$, and hence lead to a smooth quotient space $S^{5}/\Gamma$ are listed in Wolf:1974 . The abelian $\Gamma$’s are $\mathbb{Z}_{n}$ and $\mathbb{Z}_{n}\times\mathbb{Z}_{m}$. There is also a non-abelian group whose generators are subject to relations $A^{m}=B^{3n}=\mathbb{I}$ and $BAB^{-1}=A^{r}$, and in terms of SO$(6)$ matrices they are
@@
The integers $m,n,r,s$ are subject to following conditions: $(3(r-1)n,m)=(s,n)=1$ and $r\not\equiv r^{3}\equiv 1\mod m$. This subgroup was utilized in Acharya:1998db to construct conical singularities for D3333-brane probes and discussed recently in Tsimpis:2022orc . In this paper we will only analyze the $\Gamma=\mathbb{Z}_{k}$ case of $S^{5}/\Gamma$ (equivalent to a Lens space), and the non-abelian cases will be discussed in WIP . The integer homology groups of $S^{5}/\mathbb{Z}_{k}$ differ from those of $S^{5}$ only in dimension $1$ and $H_{1}(S^{5}/\mathbb{Z}_{k},\mathbb{Z})=\mathbb{Z}_{k}$.
@@
$\begin{array}[]{|c|c|c|}\hline\cr X&S^{3}/\mathbb{Z}_{k}&S^{3}/\text{Dic}_{k}\\ \hline\cr H_{0}(X)&\mathbb{Z}&\mathbb{Z}\\ \hline\cr H_{1}(X)&\mathbb{Z}_{k}&\begin{cases}\mathbb{Z}_{2}\times\mathbb{Z}_{2},&k\;\text{even}\\ \mathbb{Z}_{4},&k\;\text{odd}\end{cases}\\ \hline\cr H_{2}(X)&0&0\\ \hline\cr H_{3}(X)&\mathbb{Z}&\mathbb{Z}\\ \hline\cr\end{array}$
@@
$\begin{array}[]{|c|c|c|c|}\hline\cr X\times Y&S^{3}/\mathbb{Z}_{k}\times S^{2}&S^{3}/\text{Dic}_{k}\times S^{2}&\text{Betti number}\\ \hline\cr H_{0}(X\times Y)&\mathbb{Z}&\mathbb{Z}&b_{0}=1\\ \hline\cr H_{1}(X\times Y)&\mathbb{Z}_{k}&\begin{cases}\mathbb{Z}_{2}\times\mathbb{Z}_{2},&k\;\text{even}\\ \mathbb{Z}_{4},&k\;\text{odd}\end{cases}&b_{1}=0\\ \hline\cr H_{2}(X\times Y)&\mathbb{Z}&\mathbb{Z}&b_{2}=1\\ \hline\cr H_{3}(X\times Y)&\mathbb{Z}\oplus\mathbb{Z}_{k}&\begin{cases}(\mathbb{Z}_{2}\times\mathbb{Z}_{2})\oplus\mathbb{Z},&k\;\text{even}\\ \mathbb{Z}_{4}\oplus\mathbb{Z},&k\;\text{odd}\end{cases}&b_{3}=1\\ \hline\cr H_{4}(X\times Y)&0&0&b_{4}=0\\ \hline\cr H_{5}(X\times Y)&\mathbb{Z}&\mathbb{Z}&b_{5}=1\\ \hline\cr\end{array}$

### v3

# Berkeley, CA 94720-8162, USA

@@ Abstract
The Witten index of the $(2,0)$-theory compactified on spaces of the form $S^{3}/\Gamma\times S^{2}$, with a freely acting group $\Gamma$, and with external string sources implemented via timelike surface operator insertions, is expressed in terms of Ray-Singer torsion of $S^{3}/\Gamma$ and characters of irreducible representations of $\Gamma$. We compute it explicitly for the Dicyclic groups $\Gamma=\text{Dic}_{k}$. The torsion and characters are generally irrational numbers, but they nicely combine to an integer index. Alternatively, the Witten index can be computed from Chern-Simons theory on $S^{2}$, and Ray-Singer torsion on $S^{3}/\text{Dic}_{k}$ is thus computable from Chern-Simons theory. The matching of the Witten index calculated by these dual approaches reveals new details about the partition function of the $(2,0)$-theory with surface operators.
@@ 1 Introduction
Our goal is to describe a connection between analytic (Ray-Singer) torsion of certain free orbifolds of $S^{3}$ and elements of the Verlinde matrix of $\text{SU}(2)$ Chern-Simons theory Verlinde:1988sn ; Witten:1988hf . The connection is revealed by calculating the dimension of the Hilbert space of a certain topological compactification of the 6d $(2,0)$-theory Witten:1995zh in two different ways. We consider the $(2,0)$-theory associated with the Lie group SU(2)(2) ($A_{1}$) (i.e., the low energy limit of the worldvolume theory of two coincident M5 branes after the center of mass is decoupled Strominger:1995ac ) along a Lorentzian manifold $\mathbb{R}\times X_{5}$, with a topological twist to preserve supersymmetry, where $X_{5}$ is a generic 5-dimensional smooth Riemannian manifold with a positive scalar curvature. We will later set $X_{5}$ to a quotient of either $S^{5}$ or $S^{3}\times S^{2}$ by a freely acting finite group $\Gamma$. Compactifying on $S^{3}/\Gamma\times S^{2}$ leads to $\mathcal{N}=2$ supersymmetry in 0+1d, while compactifying on $S^{5}/\Gamma$ leads to $\mathcal{N}=1$. Our main result is (88), which is an expression for the Witten index of the theory on $S^{3}/\Gamma\times S^{2}$, with $\Gamma$ set to the Dicyclic group $\text{Dic}_{k}$ and with $\mathfrak{n}$} external string-like sources introduced via surface operators.
There is an ongoing effort to propose a fundamental formulation of the $(2,0)$-theory, and we list several approaches that have been suggested. (This list is by no means comprehensive and extensive reviews of the $(2,0)$-theory can be found in Moore2012 and Heckman:2018jxk .) The matrix model formulation of the $(2,0)$-theory on spaces with an $\mathbb{R}^{1,1}$ factor, based on instanton quantum mechanics, was proposed in Aharony:1997th , following an approach closely related to the BFSS matrix model Banks:1996vh . Another approach, known as deconstruction, was introduced in Arkani-Hamed:2001wsh . The conjecture is that in the large $N$ limit, 4d $\mathcal{N}=2$ quiver gauge theories on the Higgs branch reproduce the $(2,0)$-theory. A third proposal involves studying the $(2,0)$-theory compactified on an $S^{1}$ circle as a 5d SYM theory Douglas:2010iu ; Lambert:2010iw ; Lambert:2012qy . This idea has its origins in observations made in Seiberg:1997ax . Another approach involves the use of 3-algebras to obtain an action for multiple M5 branes, as discussed in Lambert:2010wm . This approach was motivated by the success of BLG models Bagger:2006sk ; Gustavsson:2007vu for multiple M2 branes. Finally, we mention a twistor approach to non-abelian higher gauge fields introduced in Saemann:2011nb .
Independently of the full UV description of the $(2,0)$-theory, the fundamental observation Seiberg:1997ax that 5d SYM describes the low-energy limit of the 6d theory compactified on $S^{1}$, was employed to extract remarkable exact results, using supersymmetric localization Kallen:2012cs ; Kim:2012ava ; Kim:2013nva , and it will be our main tool as well.
@@
Ccompactification on three manifolds Dimofte:2011ju results in 3d $\mathcal{N}=2$ theories referred to as Class $\mathcal{R}$. These theories exhibit a rich duality structure including mirror symmetry. In Gukov:2015sna compactification on three dimensional Seifert manifolds was studied and the resulting twisted partition function was connected with Chern-Simons theory on the base of the manifold. The last compactification we would like to mention, in the increasing dimension of compactified space, is the compactification on four manifolds $S^{3}\times S^{1}$. This results in a zero area limit of 2d q-deformed SYM theory Gadde:2009kb , where the q-deformation introduces a quantum group structure for the gauge group. This theory can actually be used to compute superconformal indices of $\text{Class}\;\mathcal{S}$ theories Gadde:2010te ; Gadde:2011ik .
In this paper, we compute the Witten index Witten:1982df of the space of ground states of the 6d $\mathcal{N}=(2,0)$ theory with SU(2)(2) gauge group, compactified on a few particular examples of $X_{5}$, with a supersymmetry-preserving R-symmetry twist. We carry out the calculations by compactifying the time direction into a Euclidean circle $S^{1}$ and shrinking the circle to small radius. In this limit Seiberg:1997ax , the 6d theory is described by an effective 5d SYM with gauge group SU(2)(2), and one finds that the circle’s radius is proportional to a square of the gauge coupling constant of the resulting 5d SYM theory. The R-symmetry twist becomes a topological twist. Hence, we can compute the Witten index by calculating the partition function of the 5d SYM theory on $X_{5}$. As explained in Aharony:1998qu ; Henningson:2010rc ; Tachikawa:2013hya , additional discrete data in the form of an element ${\mathfrak{v}}\in H^{3}(X_{5},\mathbb{Z}_{2})$ needs to be specified in order to uniquely specify what is meant by “the Hilbert space” of the $(2,0)$-theory on $X_{5}$, and in 5d SYM $\mathfrak{v}$} specifies which $\text{SU}(2)/\mathbb{Z}_{2}$ gauge bundles are to be included in the path integral.
@@
Similarly to what was originally found for 4d SYM Witten:1994cg ; Vafa:1994tf , the partition function of the 5d SYM theory and the correlation functions of SUSY Wilson loops are independent of the coupling constant and the metric because of the topological twist, which in 5d inserts a background gauge field for the $\mathfrak{so}(5)_{R}$ R-symmetry algebra that matches the geometrical spin connection of $X_{5}$. It was shown in Bershadsky:1995qy that the worldvolume theories that describe the dynamics of branes wrapping analytic submanifolds of a Calabi-Yau space naturally become topologically twisted. In our case, we can view the topological twist as the result of wrapping M5-branes on the zero section of the 10d space $T^{\star}X_{5}$ (the cotangent space of $X_{5}$). The spatial slices of the brane will reside along $X_{5}\subset T^{\star}X_{5}$, and the respective topological twist identifies the $\mathfrak{so}(5)$ holonomy algebra with the $\mathfrak{so}(5)_{R}$ R-symmetry algebra; both are subalgebras of the real superalgebra $\mathfrak{osp}(8^{*}|2)$ symmetry of the 6d $\mathcal{N}=(2,0)$ theory. This particular topological twist of the 5d SYM theory complexifies the gauge field and also results in an additional 5d Chern Simons theory with Grassmann odd sections. This twist was discussed first in Geyer:2002gy and then in Bak:2015hba ; Bak:2017jwd .
@@
In what follows, we choose $X_{5}$ to be $S^{3}/\Gamma\times S^{2}$. Here, $\Gamma$ is one of the finite subgroups of SU(2)(2) of A or D type, which are constructed by the ADE McKay correspondence mckay1981graphs . For A-type, $\Gamma$ is a cyclic group $\mathbb{Z}_{k}$, and $S^{3}/\Gamma$ is a Lens space denoted by L(k,1)L(k,1). [^1] In general one can add integers $l_{1},\dots,l_{n}$ coprime with $k$, where $n$ is the complex dimension of orbifold but in this paper we set them all to 1. For D-type, $\Gamma$ is a dicyclic group $\text{Dic}_{k}$ of order $4k$, and the resulting orbifold is referred to as a prism manifold. We will also briefly discuss $S^{5}/\Gamma$ with $\Gamma$ a cyclic group $\mathbb{Z}_{k}$ of order $k$. We review these geometries and possible extensions in detail in the first part of this paper.
We are able to compute the partition function on these geometries by applying supersymmetric localization methods. The localization equations result in BPS solutions that are given by flat connections. The $X_{5}$ spaces that we consider will have nontrivial fundamental groups and consequently, there will be nontrivial flat connections. After the path integral localization, we will need to compute functional determinants of twisted Hodge Laplacians on orbifolds. We will find that we can take a computationally simpler approach by using the Ray-Singer torsion which allows us to extract the answer by simply studying the group theoretic properties of representations of $\Gamma$.
@@
Moreover, we can compute the Witten index in a completely different way by exploiting the fact that $S^{3}/\text{Dic}_{k}$ is a circle fibration. This leads to a linear combination of dimensions of Hilbert spaces of Chern-Simons theory with quarks on $S^{2}$, and various gauge groups and levels. We will argue that the agreement of the two methods explains the simplicity of the Ray-Singer torsion of $S^{3}/\text{Dic}_{k}$, and expresses it in terms of Verlinde matrix Verlinde:1988sn ; Witten:1988hf elements.
[…]
@@ 2 Geometry of Spherical Orbifolds
We begin with a discussion of the various manifolds $X_{5}$ on which we will study the (twisted) $A_{1}$ 6d $\mathcal{N}=(2,0)$ theory, and we review some important facts about their homology groups. Recall that after compactification on a small $S^{1}$, we get a twisted version of 5d Super Yang-Mills theory with gauge group $G=\text{SU}(2)$ on the manifold $X_{5}$, which is a topological field theory that we will review below. To compute the partition function for a topological field theory that arises in this compactification, we will need to add the contributions coming from nontrivial flat connections of the SU(2)(2) gauge field. To get a nontrivial result, we will need nontrivial flat connections, which require nontrivial elements of the fundamental group $\pi_{1}(X_{5})$. Therefore, we want to pick manifolds $X_{5}$ that have a nontrivial fundamental group $\pi_{1}(X_{5})$. We will need to calculate the Ray-Singer torsion of the manifolds and to obtain non-vanishing results we will need to address the issue of fermionic zero modes. Ideal candidates for $X_{5}$ are quotients of spheres or products of spheres by a finite subgroup $\Gamma$ of the spin group.
@@
[…]
@@
Next, we consider the dicyclic group $\Gamma=\text{Dic}_{k}\rightarrow\text{SU}(2)$, with k>1k>1222Dic1 is isomorphic to $\mathbb{Z}_{4}$, which is already covered by the previous case.. The dicyclic group is a non-abelian group that is an extension of the cyclic group of order $2k$. It is generated by two elements $(r,s)$ that geometrically act as rotation and inversion respectively, and they satisfy
$r^{2k}=1,\quad s^{2}=r^{k},\quad s^{-1}rs=r^{-1}.$ (4)
We construct irreducible representations $\text{Dic}_{k}\rightarrow\text{SU}(2)$ (see also previous work GanorJu and the follow-ups Ju:2023umb ; Albrychiewicz:2024fkr that used a similar construction), which satisfy these relations for each holonomy and are given by two-dimensional matrices with unit determinant. From (4) one observes that there are abelian and non-abelian representations, which makes the manifold quotient by the dicyclic group more interesting than a quotient by a cyclic group only. We analyze these representations separately for even and odd $k$.
@@
For any even $k$, there are four one-dimensional representations, listed in Table 1 together with the notation we use to refer to them (${\mathbf{1}_{{{1}}}},\dots,{\mathbf{1}_{{{4}}}}$). Associated to them are four inequivalent embeddings of $\text{Dic}_{k}$ in SU(2)(2):
@@
where $m=1,2,\dots,k-1$, and we refer to the representation associated with a particular $m$ as $\mathbf{2}_{{m}}$}. For odd $m$, both matrices have determinant $1$, and the representation forms an embedding of Dick in SU(2)(2).
For odd $k$, the abelian irreducible representations are listed in Table 2, and the inequivalent embeddings $\text{Dic}_{k}\rightarrow\text{SU}(2)$ are listed in Table 3; there are $3$ of them. As for the non-abelian representations, the only difference from (6) is, that for odd $k$, $m$ now runs up to k2k-2.
@@
$\operatorname{Sym}^{2}{\mathbf{2}_{{m}}}=\left({\mathbf{2}_{{m}}}\otimes{\mathbf{2}_{{m}}}\right)_{S}\cong\left\{\begin{array}[]{ll}{\mathbf{2}_{{2m}}}\oplus{\mathbf{1}_{{{1}}}},&\;\text{for $m\in 2\mathbb{Z}$;}\\ {\mathbf{2}_{{2m}}}\oplus{\mathbf{1}_{{{2}}}},&\;\text{for $m\in 2\mathbb{Z}+1$.}\end{array}\right$. (7)
For odd $1\leq m\leq k-1$ the representation $\mathbf{2}_{{m}}$} is (pseudoreal) complex, while for even $m$ a reality condition can be imposed. Thus, imposing a reality condition on (7) we obtain the real $3$-dimensional representation of $\text{Dic}_{k}$ that corresponds to mapping an element $g\in\text{Dic}_{k}$ to an $\text{SU}(2)$ matrix by the representation $\mathbf{2}_{{m}}$} and then mapping the $\text{SU}(2)$ matrix to $SO(3)$ via the adjoint representation of $\text{SU}(2)$.
[…]
@@
[…]
Table 3: For odd $k$, embeddings of Dick in SU(2)(2) are constructed as direct sums of pairs of abelian representations from Table 2.
@@
We begin with the case of $S^{5}$ which has a spin isometry group Spin(6)(6), which is isomorphic to SU(4)(4). The various possible choices of $\Gamma$ that act freely on $S^{5}$, and hence lead to a smooth quotient space $S^{5}/\Gamma$ are listed in Wolf:1974 . The abelian $\Gamma$’s are $\mathbb{Z}_{n}$ and $\mathbb{Z}_{n}\times\mathbb{Z}_{m}$. There is also a non-abelian group whose generators are subject to relations $A^{m}=B^{3n}=\mathbb{I}$ and $BAB^{-1}=A^{r}$, and in terms of SO(6)(6) matrices they are
@@
The integers $m,n,r,s$ are subject to following conditions: (3(r1)n,m)=(s,n)=1(3(r-1)n,m)=(s,n)=1 and $r\not\equiv r^{3}\equiv 1\mod m$. This subgroup was utilized in Acharya:1998db to construct conical singularities for D33-brane probes and discussed recently in Tsimpis:2022orc . In this paper we will only analyze the $\Gamma=\mathbb{Z}_{k}$ case of $S^{5}/\Gamma$ (equivalent to a Lens space), and the non-abelian cases will be discussed in WIP . The integer homology groups of $S^{5}/\mathbb{Z}_{k}$ differ from those of $S^{5}$ only in dimension $1$ and $H_{1}(S^{5}/\mathbb{Z}_{k},\mathbb{Z})=\mathbb{Z}_{k}$.
@@
[…]
@@
[…]